Grade 12Physical Chemistry

If the no. of moles of free so3,
H2SO4 and H2O be x,y,z respectively in 118% H2SO4 labelled oleum,the value of(x+y+z) is

CHARISHMA
11 Years agoGrade 12

1 Answer

Ramreddy
11 Years ago
118% oleum can be preaperd by adding 10g H2SO4 and 40g SO3 and 50g water.
So, x+y+z= 10/98 + 40/80 + 50/18 = 3.38
….….….….….….….….................................................................